OpenStudy (anonymous):

What is the equation of the line in slope–intercept form that passes through the point (0, 3) and is perpendicular to the line y = 2x – 6?

OpenStudy (anonymous):

Remember to find the slope of the other line first. Since it's perpendicular the slope of the first line times the slope of the second line will equal -1. So what times 2 will get you negative 1?

OpenStudy (anonymous):

You can set this up easily by doing this: \[2x=-1\]

OpenStudy (anonymous):

im confused atm

OpenStudy (anonymous):

Where did you get confused?

OpenStudy (anonymous):

so what am i solving for first

OpenStudy (anonymous):

y=2x-6 ?

OpenStudy (anonymous):

Nope you don't need any part of that formula except for the slope of that formula. Which that slope is equal to 2 right?

OpenStudy (anonymous):

\[2x+y-3=0\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright since that slope is perpendicular to the other slope. You just need to multiple 2 times some number to get negative 1. So 2x=-1 what do you get for x? -1/2 right?

OpenStudy (anonymous):

\[x-2y-6=0\]

OpenStudy (anonymous):

okay i gotcha so far

OpenStudy (anonymous):

is the answe \[x-2y-6=0 ?\]

OpenStudy (anonymous):

y=-1/2x+3 ?

OpenStudy (anonymous):

Alright that -1/2 is the new slope of the line. So put it in slope intercept form you have y=-1/2x+b correct? now all you have to do is plug in 0 for x and 3 for y so you get 3=0+b meaning your b value is 3 So the answer to the problem is Y=-1/2x+3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Glad to help :)